9. The Ott Group, Inc., has identified the following two mutually exclusive projects: Year 0 Cash Flow (S) - $12,500 4,000 5,please show full steps of the solution for full credit. not excel solutions please. thank you!!!
Answer #1

a)

Project S:

IRR is the rate of return that makes initial investment equal to present value of cash inflows

12500 = 4000 / (1 + r)1 + 5000 / (1 + r)2 + 6000 / (1 + r)3 + 1000 / (1 + r)4

Using trial and error method, i.e., after trying various values for R, let try R as 11.85%

12500 = 4000 / (1 + 0.1185)1 + 5000 / (1 + 0.1185)2 + 6000 / (1 + 0.1185)3 + 1000 / (1 + 0.1185)4

12500 = 12500

Therefore, IRR of project S is 11.85%

Project L:

IRR is the rate of return that makes initial investment equal to present value of cash inflows

12500 = 1000 / (1 + r)1 + 6000 / (1 + r)2 + 5000 / (1 + r)3 + 4000 / (1 + r)4

Using trial and error method, i.e., after trying various values for R, let try R as 9.53%

12500 = 1000 / (1 + 0.0953)1 + 6000 / (1 + 0.0953)2 + 5000 / (1 + 0.0953)3 + 4000 / (1 + 0.0953)4

12500 = 12500

Therefore, IRR of project L is 9.53%

Project S should be accepted as it has the higher IRR

b)

Project S:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -12500 + 4000 / (1 + 0.11)1 + 5000 / (1 + 0.11)2 + 6000 / (1 + 0.11)3 + 1000 / (1 + 0.11)4

NPV = $207.60

Project L:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -12500 + 1000 / (1 + 0.11)1 + 6000 / (1 + 0.11)2 + 5000 / (1 + 0.11)3 + 4000 / (1 + 0.11)4

NPV = -438.48

Project S should be accepted as it has a positive NPV